is the solution which contains 0.3 m of hno3 3. is a solution containing 2.22 g of x (oh) 2 250 cm cube of solution. during titration 19.0 cm cube of a completely neutralise 25.0 cm cube of bcalculate the molar mass of x (oh)2
1
Expert's answer
2019-11-15T07:06:48-0500

M(X(OH)2)=m(X(OH)2)*2*1000/(V(HNO3)*C(HNO3))=2.22*2*100/(19*0.3)=78 (g/mol)

Answer: 78 g/mol.